OpenStudy (anonymous):

Matrix question-Tell whether the product AB is defined for A4x5 and B5x3. If yes, give the product's dimensions as the following example: 2x3

OpenStudy (anonymous):

As long as the two inner numbers, the 5 (column) of A and the 5 (row) of B, are the same then the product of the two matrices is defined. The dimensions will be the two outer numbers, the 4 (row) of A and the 3 (column of B), so 4x3 is the resulting dimension.

OpenStudy (anonymous):

mbradar2 is right !!

OpenStudy (maheshmeghwal9):

\[4 \times \cancel 5<--> \cancel5 \times 3.\]So matrices can be multiplied. & Order of product is \[4 \times 3.\]